An object has a kinetic energy of 175 J and a momentum of magnitude 25.0 kg m/s. Find the speed of the object.
1
Expert's answer
2020-05-13T09:57:27-0400

Let denotes the kinetic energy of the object is "K" , momentum is "p" and mass is "m."

We know that,

"K=\frac{1}{2}mv^2 \hspace{1cm}(1)\\
p=mv\hspace{1cm} (2)"

We have given that,


"K=175J\\
p=25.0kg\:ms^{-1}"

Now. divide the equation "(1)" by "(2)" we get,


"\frac{K}{p}=\frac{\frac{1}{2}mv^2}{mv}\\
\implies \frac{2K}{p}=v \hspace{1cm}(3)"

Thus, from equation "(3)" we get speed of the object is

"v=\frac{2\times 175}{25}=14ms^{-1}"

Hence, we are done.